# Computer-Assisted Proofs of Solitons in Bose-Einstein Condensates

### v1

@@ 2.1 Solving for The First Order Coefficients: The Stable Bundle
Since we plan to use computers for our calculations, we need to determine what mathematical objects can be implemented computationally. A mathematical object is called computable if it can be explicitly implemented in a computational program, meaning it can be constructed or evaluated through a finite sequence of well-defined computational steps.
Computable expressions can be rigorously evaluated using interval arithmetic, where numbers are represented as intervals instead of single floating-point values. Arithmetic operations such as addition, multiplication, and division are performed directly on these intervals. This approach propagates numerical uncertainty throughout the computation, producing intervals that rigorously enclose the true result.
@@ 2.2 Solving for Higher Order Coefficients
We can only guarantee that the output of a parameterization coming from Theorem 11 is included in a ball of some radius. However, this fact does not represent a computational limitation. We are now ready to solve the boundary value problem (7).
3 Solving the Boundary Value Problem
Having detailed in the previous section the computation of a parameterization $W$ of the stable manifold associated with the periodic orbit $\gamma$, we can now reformulate the boundary value problem (7) as follows
@@ 3 Solving the Boundary Value Problem
We proceed with the definition of the validation map for the soliton boundary value problem. Notably, we employ the same notation as in previous contexts to highlight the structural parallels in our approach. As derived in Section 2, let $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ be a parameterization of the local stable manifold $W^{s}_{\rm loc}(\gamma)$ associated with the periodic orbit $\gamma$. The validation map is defined as follows:
@@
In the following section we provide an example of how to use Theorem 15 to validate a numerical approximate soliton solution.
4 Constructive proofs of existence of solitons
We now present our computer-assisted method for proving the existence of soliton solutions to equation (2) for different parameter sets $\{a,b,c\}$. Applying Theorems 7, 11, and 15, we demonstrate that true soliton solutions exist near numerically approximated ones.
The computer-assisted component of this work consists of verifying the hypotheses of the theorems through computational tools. This requires implementing all computable expressions introduced in previous sections, including operators, derivatives, and bounds. We use the Julia programming language and the RadiiPolynomial.jl package to manipulate sequences and projection operators. To maintain mathematical rigor, we employ interval arithmetic via the IntervalArithmetic.jl package. Each implementation is designed to work directly with interval arithmetic.
To prove inequalities, we evaluate computable bounds using interval arithmetic, which outputs an interval containing the true value. The right endpoint of this interval serves as a rigorous upper bound. In this way, we can verify all computable bounds in the hypotheses of Theorems 711, and 15 using a computer. As a first example, we restate Theorem 1 in more detail and provide a computer-assisted proof. As our first example, we restate Theorem 1 in more detail and provide a computer-assisted proof.
@@ 4 Constructive proofs of existence of solitons
$\|u-\bar{u}\|_{\infty}\leq 8.617584260554394\times 10^{-6},$
where $\bar{u}$ is a numerical approximation of the solution illustrated in Figure 1. The data of the approximate solution can be found in [38].
@@
We choose truncation modes $M_{\scriptscriptstyle{\mathsf{F}}}$, $N_{\scriptscriptstyle{\mathsf{T}}}$, and $M_{\scriptscriptstyle{\mathsf{C}}}$ to capture the nonlinear behavior of the equation, so that the final coefficients of each sequence have decayed to the level of rounding error in double precision.
For this particular set of parameters, we have obtained $\bar{x}_{\scriptscriptstyle{\mathsf{F}}}\in X_{\scriptscriptstyle{\mathsf{F}}}$, $\bar{x}_{\scriptscriptstyle{\mathsf{TF}}}\in X_{\scriptscriptstyle{\mathsf{TF}}}$, and $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}\in X_{\scriptscriptstyle{\mathsf{C}}}$ satisfying (70) for $M_{\scriptscriptstyle{\mathsf{F}}}=32$, $N_{\scriptscriptstyle{\mathsf{T}}}=32$ and $M_{\scriptscriptstyle{\mathsf{C}}}=48$. Together with the following residual conditions
@@
For the Fourier expansion, we use the same truncation mode for both the bundle and the manifold, as we can always take the maximum between the two when needed. The coefficients used in these computations are available in the GitHub repository associated with this paper [38].
The zero-finding problems (27), (44), and (61) define well-conditioned maps that can be used with root-finding algorithms to refine numerical approximations. In practice, after obtaining an approximate solution, we apply Newton’s method to these maps to improve the accuracy of our numerical approximations. The required level of accuracy varies depending on the proof. In some cases, smaller residuals are necessary, while in others, lower accuracy suffices.
Our proof consists of three stages: the proof of the tangent bundle, the proof of the stable manifold attached to the periodic orbit and the validation of the solution of the soliton boundary value problem. We begin by obtaining a solution to the bundle problem. Our numerical approximation $\bar{x}_{\scriptscriptstyle{\mathsf{F}}}\in X_{\scriptscriptstyle{\mathsf{F}}}$ of the bundle problem satisfies that $\bar{\lambda}<0$ and the coefficients of $\bar{v}$ satisfy the symmetry condition
@@
Applying Theorem 7, we obtain a unique zero of the validation map (27),
@@
within the ball $B_{X_{\scriptscriptstyle{\mathsf{F}}}}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}},r_{\scriptscriptstyle{\mathsf{F}}})$ with radius $r_{\scriptscriptstyle{\mathsf{F}}}=4.122891017172993\times 10^{-13}$. This zero corresponds to a solution
$(\lambda,v^{(1)},v^{(2)},0,0)\in\mathbb{R}\times S^{4}_{\scriptscriptstyle{\mathsf{F}}},$
which satisfies the sequence equation (23) and represents a real solution of (11). With this result, we can now evaluate the validation map and its derivative for the parameterization of the stable manifold attached to the periodic orbit described by (44).
@@
Therefore, we use Theorem 11 to validate our paramaterization of the manifold $x_{\scriptscriptstyle{\mathsf{TF}}}\in X_{\scriptscriptstyle{\mathsf{TF}}}$ with a radius $r_{\scriptscriptstyle{\mathsf{TF}}}=1.5204458252945915\times 10^{-7}$. The resulting stable manifold is represented in orange in Figure 1.
Once we obtain the stable manifold via the $W$ parameterization, we compute a numerical approximation to the corresponding boundary value problem. In this case, setting $\theta=1$ and $\mathsf{L}=1+2\pi$ in (57), we obtain
@@
Next, in a manner analogous to the previous stages, we fix a norm weight $\omega=1.05$ in (15) and compute the following bounds:
@@
for $r_{\scriptscriptstyle{\mathsf{C}}}^{*}=10^{-2}$. Notice that $|\bar{\sigma}+r_{\scriptscriptstyle{\mathsf{C}}}^{*}|<1$, as required by Lemma 14. We then apply Theorem 15 to validate the solution of the boundary value problem $x_{\scriptscriptstyle{\mathsf{C}}}\in X_{\scriptscriptstyle{\mathsf{C}}}$. This yields a solution $x=(\sigma,s)\in X_{\scriptscriptstyle{\mathsf{C}}}$ such that
@@
which solves the boundary value problem (7) in the ball centered at $\bar{x}=(\bar{\sigma},\bar{s})$ with radius
$r_{\scriptscriptstyle{\mathsf{C}}}=8.617584260554394\times 10^{-6}.$
The solution of the boundary value problem is depicted in blue in Figure 1. Moreover, the periodic potential $V(x)=\cos(2x)$ satisfies the initial value problem $u_{3}^{\prime\prime}=-4u_{3}$ with initial conditions $(u_{3}(0),u_{3}^{\prime}(0))=(1,0)$. And therefore if $\mathbf{u}=(u_{1},\dots,u_{4})\in\mathbb{R}^{4}$ is the solution in $[0,\infty)$ then the even extension of $u_{1}$ is a solution of equation (2). The proven soliton solution satisfies
$\|\bar{s}^{(i)}-s^{(i)}\|_{\infty}=\sup_{t\in[-1,1]}\|\bar{s}^{(i)}(t)-s^{(i)}(t)\|_{\scriptscriptstyle{\mathsf{C}}}\leq r_{\scriptscriptstyle{\mathsf{C}}}=8.617584260554394\times 10^{-6},\quad i=1,2,3,4.$
The implementations of the proofs with interval arithmetic of all the inequalities above are included in [38]. ∎
Our method can be easily adapted to different sets of parameters. For example, we provide existence proofs for the solitons presented in [10], including the one- and two-hump solutions illustrated in Figure 2.
Figure 2: Proofs for the solitons originally presented in [10]. Projected boundary problem (blue), stable manifold (orange), and even extension (black).

### v2

@@ 2.1 Solving for The First Order Coefficients: The Stable Bundle
Since we plan to use computers for our calculations, we need to determine what mathematical objects can be implemented computationally. A mathematical object is called computable if it can be explicitly implemented in a computational program, meaning it can be constructed or evaluated through a finite sequence of well-defined computational steps. Computable expressions can be rigorously evaluated using interval arithmetic, where numbers are represented as intervals instead of single floating-point values. Arithmetic operations such as addition, multiplication, and division are performed directly on these intervals. This approach propagates numerical uncertainty throughout the computation, producing intervals that rigorously enclose the true result (e.g., see [31, 26]).
@@ 2.2 Solving for Higher Order Coefficients
We can only guarantee that the output of a parameterization coming from Theorem 11 is included in a ball of some radius. However, this fact does not represent a computational limitation. We are now ready to solve the boundary-value problem (7).
3 Solving the Boundary-Value Problem
Having detailed in the previous section the computation of a parameterization $W$ of the stable manifold associated with the periodic orbit $\gamma$, we can now reformulate the boundary-value problem (7) as follows
@@ 3 Solving the Boundary-Value Problem
We proceed with the definition of the validation map for the soliton boundary-value problem. Notably, we employ the same notation as in previous contexts to highlight the structural parallels in our approach. As derived in Section 2, let $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ be a parameterization of the local stable manifold $W^{s}_{\rm loc}(\gamma)$ associated with the periodic orbit $\gamma$. The validation map is defined as follows:
@@
We are now ready to validate numerically approximated soliton solutions.
4 Examples of Constructive Proofs of Existence of Gap Solitons
In this section, we present examples of our computer-assisted method for proving the existence of soliton solutions to equation (2). Using Theorems 7, 11, and 15, we demonstrate that true soliton solutions exist near numerically approximated ones.
Given a numerical approximation to a soliton solution, our approach requires implementing all relevant expressions—operators, derivatives, and bounds—introduced in earlier sections. The computer-assisted component comes in the form of rigorously evaluating these quantities using interval arithmetic. For our examples, we use the Julia programming language and we use interval arithmetic through the IntervalArithmetic.jl Julia package [32]. Our code implementation is available in the repository associated with this paper [33].
To verify all bounds in the hypotheses of Theorems 7, 11, and 15, we evaluate each computable quantity using intervals rather than individual floating-point numbers. This guarantees that the true value lies within the resulting interval. The right endpoint provides a rigorous upper bound for the quantity being estimated. As a first example, we restate Theorem 1 and provide a computer-assisted proof.
@@ 4 Examples of Constructive Proofs of Existence of Gap Solitons
$\|u-\bar{u}\|_{\infty}\leq 8.617584260554394\cdot 10^{-6},$
where $\bar{u}$ is a numerical approximation of the solution illustrated in Figure 1.
@@
We choose truncation modes $M_{\scriptscriptstyle{\mathsf{F}}}$, $N_{\scriptscriptstyle{\mathsf{T}}}$, and $M_{\scriptscriptstyle{\mathsf{C}}}$ to capture the nonlinear behavior of the equation, so that the final coefficients of each sequence have decayed to the level of rounding error in double precision. For this particular set of parameters, we have obtained $\bar{x}_{\scriptscriptstyle{\mathsf{F}}}\in X_{\scriptscriptstyle{\mathsf{F}}}$, $\bar{x}_{\scriptscriptstyle{\mathsf{TF}}}\in X_{\scriptscriptstyle{\mathsf{TF}}}$, and $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}\in X_{\scriptscriptstyle{\mathsf{C}}}$ satisfying (70) for $M_{\scriptscriptstyle{\mathsf{F}}}=32$, $N_{\scriptscriptstyle{\mathsf{T}}}=32$ and $M_{\scriptscriptstyle{\mathsf{C}}}=48$. Together with the following residual conditions
@@
For the Fourier expansion, we use the same truncation mode for both the bundle and the manifold, as we can always take the maximum between the two when needed. The coefficients used in these computations are available in the repository associated with this paper [33]. In our implementation, we use the RadiiPolynomial.jl Julia package [32] to easily manipulate sequences.
The required level of numerical accuracy depends on the solution we aim to validate. However, as a byproduct of our approach, the zero-finding problems (27), (44), and (61) define well-conditioned maps that can be used with root-finding algorithms to refine our numerical approximations. In practice, after computing an initial approximation, we apply Newton’s method to these maps to improve accuracy to the required level.
Our proof has three stages. We begin with the bundle problem, whose solution provides the input needed to evaluate the validation map for the stable manifold attached to the periodic orbit defined by (27). Next, we construct and validate the parameterization of this manifold. From this, we obtain an explicit boundary condition for the boundary-value problem (7), which we then use to validate a numerical approximation of the soliton solution.
For the first stage, the numerical approximation of the bundle problem $\bar{x}_{\scriptscriptstyle{\mathsf{F}}}=(\bar{\lambda},\bar{v})\in X_{\scriptscriptstyle{\mathsf{F}}}$ satisfies $\bar{\lambda}<0$. Moreover, the coefficients of $\bar{v}$ satisfy the symmetry condition
@@
Note that this condition can be directly imposed and verified in the computational implementation. We fix the norm weight $\nu=1.05$ in (15) and the scaling factor $l=0.5$ in the phase condition (26). Using interval arithmetic, we validate the following bounds:
@@
as defined in (32), (33), and (41). These bounds satisfy the inequalities
$Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<1\quad\text{and}\quad Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<\frac{(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}}{2Y(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})},$
which meet the hypotheses of Theorem 7. We therefore obtain a unique zero of the validation map (27),
@@
within the ball $B_{X_{\scriptscriptstyle{\mathsf{F}}}}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}},r_{\scriptscriptstyle{\mathsf{F}}})$ of radius $r_{\scriptscriptstyle{\mathsf{F}}}=4.122891017172993\cdot 10^{-13}$. This zero defines a solution $(\lambda,v^{(1)},v^{(2)},0,0)\in\mathbb{R}\times S^{4}_{\scriptscriptstyle{\mathsf{F}}},$ to the sequence equation (23) and corresponds to a real solution of the bundle differential equation (11).
@@
Therefore, we use Theorem 11 to validate our parameterization of the manifold $x_{\scriptscriptstyle{\mathsf{TF}}}\in X_{\scriptscriptstyle{\mathsf{TF}}}$ with a radius $r_{\scriptscriptstyle{\mathsf{TF}}}=1.5204458252945915\cdot 10^{-7}$. The resulting stable manifold is represented in orange in Figure 1.
@@
for $r_{\scriptscriptstyle{\mathsf{C}}}^{*}=10^{-2}$. Notice that $|\bar{\sigma}+r_{\scriptscriptstyle{\mathsf{C}}}^{*}|<1$, as required by Lemma 14. We then apply Theorem 15 to obtain a solution $x=(\sigma,s)\in X_{\scriptscriptstyle{\mathsf{C}}}$ such that
@@
solves the boundary-value problem (7) (after reverting the scaling in time) in the ball centered at $\bar{x}=(\bar{\sigma},\bar{s})$ with radius $r_{\scriptscriptstyle{\mathsf{C}}}=8.617584260554394\cdot 10^{-6}.$ The solution of the boundary-value problem is depicted in blue in Figure 1.
As described in the introduction, by taking the even extension of a solution to the boundary-value problem (7), we obtain a soliton solution to the Gross-Pitaevskii equation (2). The proven soliton solution satisfies
$\|\bar{s}^{(i)}-s^{(i)}\|_{\infty}=\sup_{t\in[-1,1]}\|\bar{s}^{(i)}(t)-s^{(i)}(t)\|_{\scriptscriptstyle{\mathsf{C}}}\leq r_{\scriptscriptstyle{\mathsf{C}}}=8.617584260554394\cdot 10^{-6},\quad i=1,2,3,4.$
Observe that, by (55), the error bound also applies to the portion of the solution obtained via the parameterization of the stable manifold, not just the segment in the boundary-value problem region. The proofs with interval arithmetic of all the inequalities above are included in [33]. ∎
The implementation of our method can be easily adapted to different sets of parameters. For example, we provide existence proofs for the solitons presented in [4], including the one- and two-hump solutions illustrated in Figure 2.
Figure 2: Computer-assisted proofs for numerically approximated solitons originally presented in [4]. The parameters of the equation are shown at the top of the figure. The illustration also shows the main components of our approach: the solution to the boundary-value problem (blue), the stable manifold (orange), and the even extension of the soliton (black).
@@
Author Contributions Statement
All authors equally contributed to this manuscript and should be listed in alphabetical order.
Funding
JP L. was supported by NSERC.